OpenStudy (anonymous):

Solve by the elimination method. 8x-9y=8.5, 7y-2x=-4.5

OpenStudy (anonymous):

first you want to set the up top each other and decide which variable you want to cancel out

OpenStudy (anonymous):

8x-9y=8.5 7y-2x=-4.5

OpenStudy (anonymous):

The first one?

OpenStudy (anonymous):

yes 8x-9y=8.5 -2x+7y=-4.5

OpenStudy (anonymous):

what do you do next?

OpenStudy (anonymous):

assuming we want to get rid of x we multpliy the first equation by -2 and the second equation by 8 thats the whole equation on both sides

OpenStudy (anonymous):

so it would be -16x-18y=-17, 56y-126x=-36?

OpenStudy (anonymous):

@Jasmine2012

OpenStudy (anonymous):

no it would be -16x+18y=-17 -16x-56y=-36

OpenStudy (anonymous):

then add them together to get rid of the x's

OpenStudy (anonymous):

then it would be y=-53/74

OpenStudy (anonymous):

38/53

OpenStudy (jdiegosantillan):